OpenStudy (anonymous):

the length of a rectangle is 3 more com that twice the width the perimeter is 150 and i have to find the length

zepdrix (zepdrix):

"3 more com? that twice the.." What did you mean to say? D:

OpenStudy (anonymous):

cm*

OpenStudy (anonymous):

so let L be the length of the rectangle, and W be the width. Perimeter = 2L + 2W = 150 L = (W+3)/2 Plug that equation into the perimeter equation to find your answer

OpenStudy (anonymous):

whoops, sorry. L = 2*(W+3)

OpenStudy (anonymous):

My god, I'm making mistakes everywhere... L = 2W + 3

OpenStudy (anonymous):

Sorry about that. Plug L = 2W + 3 into the perimeter equation 150 = 2L + 2W You can either find the width, and then find the length, or directly find the length by solving L = 2W + 3 for W and plugging that into he equation